# Dual-Spaces Invariance as a Universal Criterion for Identifying Multifractal Critical States

### v5

@@ III.3 Verification results
Figure 3: Mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for quasiperiodic models. (a) AAH model: the two quantities exhibit markedly different finite-size scaling behaviors away from the critical point $V=2$, while they become comparable at criticality. The data are obtained from all eigenstates for system sizes L=500, 1000, 2000L=500,\;1000,\;2000. (b) QNE model: in the regime $0<V<2$, both MIPR and MIPRm show strong fluctuations and similar scaling tendencies, consistent with critical behavior over an extended parameter range. The data are obtained from six eigenstates near $E=0.00001$ for system sizes L=500, 1000, 2000L=500,\;1000,\;2000.
@@
We perform direct numerical diagonalization for systems with $L=987$ lattice sites to obtain the eigenvalues and eigenstates of both the AAH and QNE models. The numerical results are presented in Fig. 2. As shown in panels (a1), (a2) and (b1), (b2), the critical states in both models exhibit simultaneous delocalization, multifractality, and self-similarity in both position space ($|\psi|$) and momentum space ($|\phi|$). These results confirm the theoretical assertion that the wave function of the critical state retains a certain invariance under duality transformation.
@@
Figure 3 presents a quantitative comparison between the mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for the two representative quasiperiodic models considered in this work. To reduce sample-to-sample fluctuations, the MIPR is obtained by averaging the IPR over a set of eigenstates near a fixed energy. For the AAH model, we take the number of sampled states to scale with the system size, whereas for the computationally more demanding QNE model we restrict the averaging to a small number of eigenstates.
@@
As shown in Fig. 3(a), the AAH model exhibits a pronounced asymmetry between position and momentum space away from the critical point. In the regime $V<2$, the real-space MIPR displays strong system-size dependence, consistent with extended states, while the momentum-space MIPRm approaches a size-independent value, indicative of localization in momentum space. The opposite behavior is observed for $V>2$. At the self-dual point $V=2$, corresponding to criticality, the two quantities become comparable in magnitude and exhibit similar scaling behavior, in accordance with the dual-space correspondence discussed in Sec. II.
@@
The behavior of the QNE model, shown in Fig. 3(b), is qualitatively different. For $V>2$, the eigenstates are exponentially localized in real space, leading to a weak size dependence of the MIPR, while the momentum-space MIPRm exhibits pronounced scaling with system size. In contrast, in the regime $0<V<2$, both MIPR and MIPRm display strong fluctuations and nontrivial size dependence. Rather than approaching well-defined plateaus characteristic of extended or localized phases, the two quantities exhibit comparable scaling tendencies, which is consistent with the presence of critical states over an extended parameter range.
@@
We emphasize that, due to finite-size effects and the basis dependence of the IPR, an exact equality between MIPR and MIPRm is neither expected nor required. Instead, the numerical results demonstrate that critical states are characterized by comparable scaling behavior between position and momentum space. These observations provide direct numerical support for the dual-space correspondence formulated in Sec. II.
@@ IV Experimental feasibility
The dual-space correspondence discussed in this work is particularly well suited for experimental verification in ultracold atomic systems, where both real-space and momentum-space observables can be accessed with high precision. One-dimensional quasiperiodic lattices realizing AAH-type Hamiltonians have been extensively implemented using optical lattices with superimposed incommensurate potentials g1 ; gg9 .
@@
In such setups, the real-space density profile $|\psi_{n}|^{2}$ can be directly measured via in situ fluorescence or absorption imaging, enabling the extraction of the real-space inverse participation ratio $\mathrm{IPR}$. Complementarily, the momentum-space distribution $|\phi_{k}|^{2}$ can be obtained through standard time-of-flight (TOF) measurements. After suddenly switching off the trapping and lattice potentials, the atomic cloud undergoes ballistic expansion, and in the long-TOF limit the measured density profile is proportional to the initial momentum distribution. By discretizing the TOF images into momentum bins determined by the imaging resolution, the momentum-space inverse participation ratio $\mathrm{IPR}_{m}$ can be evaluated in a directly analogous manner.
@@
Within this experimental framework, extended and localized states exhibit a pronounced asymmetry between position and momentum space, whereas critical states are expected to display comparable localization characteristics in both representations. Although finite-size effects and experimental resolution prevent an exact equality between $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$, the scaling-level correspondence identified in this work should remain observable within realistic system sizes. Our findings thus establish a practical framework for diagnosing multifractal critical states with currently available cold-atom quantum simulators.
V Summary
@@
In summary, we have investigated critical states in one-dimensional quasiperiodic systems from a dual-space perspective, focusing on the relationship between position and momentum representations. By combining analytical considerations with numerical simulations of representative models, including the AAH model and a quasiperiodic nonlinear eigenproblem, we have shown that critical states exhibit closely related localization characteristics in the two dual spaces. This behavior contrasts sharply with extended or localized states, which display strong asymmetry between real and momentum space.
@@
Beyond Lyapunov exponents, we have demonstrated that the inverse participation ratio show comparable scaling behavior in position and momentum space for critical states. Although this correspondence does not imply an exact invariance at the level of wave-function moments, it provides a robust and practical criterion for identifying multifractal criticality in finite systems.
Our results highlight the usefulness of dual-space analysis as an operational tool for diagnosing critical states in quasiperiodic lattices. An important open question is whether similar dual-space correspondence persists in higher-dimensional or genuinely disordered systems, where the structure of momentum space is more complex. Addressing this issue would further clarify the scope and general applicability of the dual-space characterization of critical states.
@@ 1. Real-space Lyapunov exponent $\gamma$
For a given energy $E$ and a fixed realization of the quasiperiodic or disordered potential, the Lyapunov exponent $\gamma(E)$ characterizes the asymptotic exponential growth rate of wave-function amplitudes along the lattice. It can be computed efficiently using the standard transfer-matrix method. Starting from an initial vector $\mathbf{v}_{0}$ (e.g., $(1,0)^{\mathrm{T}}$), we iteratively apply the site-dependent $2\times 2$ transfer matrix $\mathbf{T}_{n}(E)$,
@@
After $N$ iterations, the total transfer matrix reads $\boldsymbol{\Pi}_{N}(E)=\mathbf{T}_{N-1}\cdots\mathbf{T}_{1}\mathbf{T}_{0}$. The Lyapunov exponent is defined as
$\gamma(E)=\lim_{N\to\infty}\frac{1}{N}\ln\left\|\boldsymbol{\Pi}_{N}(E)\right\|,$ (10)
where $\|\cdot\|$ denotes a suitable vector norm. In numerical calculations, the vector is periodically renormalized to avoid overflow, and $\gamma(E)$ is extracted from the average logarithmic growth rate over a sufficiently long iteration ($N\sim 10^{5}$).
@@ 2. Momentum-space Lyapunov exponent $\gamma_{m}$
The momentum-space Lyapunov exponent $\gamma_{m}$ is computed using an analogous transfer-matrix approach formulated in momentum space. After Fourier transforming the real-space eigenvalue problem, one obtains a linear recurrence relation for the momentum-space amplitudes $\tilde{\psi}_{k}$. For quasiperiodic potentials with a single incommensurate frequency, this recurrence typically couples momentum modes forming an effective one-dimensional chain. Truncating the coupling beyond a sufficiently large cutoff leads to a finite-difference equation of the form
@@
where $Q$ denotes the characteristic momentum spacing. This equation can be recast into a $2\times 2$ transfer-matrix form,
$\begin{pmatrix}\tilde{\psi}_{k+Q}\\ \tilde{\psi}_{k}\end{pmatrix}=\mathbf{M}_{k}(E)\begin{pmatrix}\tilde{\psi}_{k}\\ \tilde{\psi}_{k-Q}\end{pmatrix}$, (12)
with the matrix elements of $\mathbf{M}_{k}(E)$ determined by $a_{k}$, $b_{k}$, and $c_{k}$. Iterating this relation while periodically renormalizing the vector yields the momentum-space Lyapunov exponent,
@@
where $\boldsymbol{\Pi}^{(m)}_{K}(E)$ denotes the product of momentum-space transfer matrices over $K$ steps.
In both representations, the Lyapunov exponent quantifies the exponential sensitivity of the wave-function amplitudes to lattice or momentum index. The difference lies solely in the space where the linear recurrence is defined and in the specific form of the corresponding transfer matrices.
Appendix B Remark: Basis dependence of the IPR versus Lyapunov exponents
It is instructive to clarify the distinct roles played by the Lyapunov exponent and the inverse participation ratio (IPR) in the characterization of localization properties. The Lyapunov exponent is defined through the asymptotic exponential growth (or decay) rate of solutions to a linear recurrence relation and is therefore an intrinsic dynamical quantity associated with the transfer-matrix structure of the system. As such, it does not depend on the choice of basis used to represent the wave function, but rather on the underlying recursive dynamics in a given representation (real or momentum space).
By contrast, the IPR explicitly depends on the expansion coefficients of a wave function in a chosen basis and quantifies the degree of sparsity or concentration of probability weight with respect to that representation. Since higher-order moments of wave-function amplitudes are not invariant under general unitary transformations, the IPR is inherently basis dependent and generally takes different values in position and momentum space.
This distinction is essential for the present work. While the dual-space vanishing of Lyapunov exponents provides a representation-independent criterion for criticality, the emergence of closely related scaling behavior of the IPR in dual spaces for critical states constitutes a nontrivial and physically meaningful observation. It reflects the fact that, despite the basis dependence of the IPR, critical states exhibit reduced asymmetry between position and momentum representations at the level of scaling properties.
@@ Appendix C Fourier transformation for the QNE model
We demonstrate that the real-space equation
@@
is transformed, under a discrete Fourier transform, into the momentum-space equation
@@
1. Preliminaries
To eliminate the nonlocal structure of $\tan x=\sin x/\cos x$, we multiply Eq. (14) by $\cos(2\pi\alpha n)$, yielding
@@ 1. Preliminaries
2. Transformation of individual terms
Using standard trigonometric identities and Fourier algebra, one finds
@@ 2. Transformation of individual terms
3. Momentum-space equation
Substituting the above relations into the Fourier transform of Eq. (16), we arrive at
@@ 3. Momentum-space equation
which can be rearranged into the compact form given in Eq. (15). This completes the derivation of the momentum-space representation of the QNE model.

### v6

@@ III.3 Verification results
Figure 3: (a) AAH model: MIPR and MIPRm exhibit distinct scaling away from the critical point $V=2$, but converge to comparable values at criticality. Results are averaged over all eigenstates for L=500, 1000, 2000L=500,\;1000,\;2000. (b) QNE model: in the regime $0<V<2$, both MIPR and MIPRm display strong fluctuations yet remain comparable, supporting criticality over a broadened parameter range. Results are obtained from six eigenstates near $E=0.00001$ for L=500, 1000, 2000L=500,\;1000,\;2000.
@@
We perform direct numerical diagonalization for systems with $L=987$ to obtain the eigenvalues and eigenstates of both the AAH and QNE models. The results are shown in Fig. 2. As seen in panels (a1), (a2) and (b1), (b2), the critical states exhibit simultaneous delocalization, multifractality, and self-similarity in both position space ($|\psi|$) and momentum space ($|\phi|$). These results provide direct numerical evidence that critical wave functions obey dual-space invariance under Fourier transformation.
@@
Figure 3 compares the mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for the two quasiperiodic models. The MIPR is computed by averaging the IPR over eigenstates within a narrow energy window to suppress sample fluctuations. For the AAH model, the number of sampled states scales with system size, whereas for the QNE model, the averaging is limited to a few eigenstates due to computational cost.
@@
As shown in Fig. 3(a), the AAH model exhibits strong asymmetry between position and momentum space away from criticality. For $V<2$, the real-space MIPR decreases with system size, consistent with extended states, while the momentum-space MIPRm remains finite, indicating localization in momentum space. The opposite behavior occurs for $V>2$. At the self-dual point $V=2$, the two quantities become comparable and follow similar scaling, reflecting dual-space invariance at criticality.
@@
The behavior of the QNE model, shown in Fig. 3(b), is qualitatively different. For $V>2$, the eigenstates are exponentially localized in real space, leading to a weak size dependence of the MIPR, while the momentum-space MIPRm exhibits pronounced system-size scaling. In contrast, for $0<V<2$, both MIPR and MIPRm show strong fluctuations and nontrivial size dependence. Rather than approaching plateaus characteristic of extended or localized phases, the two quantities exhibit comparable scaling, consistent with a critical regime.
@@
We emphasize that, due to finite-size effects and the basis dependence of the IPR, exact equality between MIPR and MIPRm is neither expected nor required. Instead, the results show that critical states are characterized by comparable scaling in position and momentum space. These findings provide direct numerical support for the dual-space correspondence.
@@ IV Experimental feasibility
The dual-space correspondence identified in this work is well suited for experimental verification in ultracold atomic systems, where both position- and momentum-space observables are directly accessible. One-dimensional quasiperiodic lattices realizing AAH-type Hamiltonians can be implemented using optical lattices with superimposed incommensurate potentials.
@@
In such setups, the real-space density distribution $|\psi_{n}|^{2}$ can be measured using in situ fluorescence or absorption imaging gg9 , allowing direct evaluation of the real-space IPR. The momentum-space distribution $|\phi_{k}|^{2}$ can be obtained via standard time-of-flight (TOF) measurements Biddle . After switching off the trapping and lattice potentials, the atomic cloud undergoes ballistic expansion, and in the long-TOF limit the measured density profile reflects the initial momentum distribution. By discretizing the TOF images according to the imaging resolution, the momentum-space inverse participation ratio $\mathrm{IPR}_{m}$ can be extracted in a directly analogous manner.
@@
Within this framework, extended and localized states exhibit pronounced asymmetry between position and momentum space, whereas critical states show comparable localization properties in both representations. Although finite system sizes and experimental resolution prevent exact quantitative agreement, the scaling-level correspondence predicted here remains observable under realistic conditions. These results demonstrate that current cold-atom experiments provide a feasible and practical route for identifying multifractal critical states through dual-space diagnostics.
V Summary and outlook
@@
In summary, we have investigated critical states in one-dimensional quasiperiodic systems from a dual-space perspective. Using analytical arguments and numerical simulations of representative models, including the AAH model and a quasiperiodic nonlinear eigenproblem, we have shown that critical states are uniquely characterized by comparable scaling in dual spaces. This property sharply distinguishes them from extended and localized states, which exhibit a pronounced asymmetry between position and momentum spaces.
@@
We have further demonstrated that the inverse participation ratio provides a simple and practical diagnostic of this behavior: critical states exhibit comparable scaling of the inverse participation ratio in position and momentum space even in finite systems. This result establishes dual-space correspondence as an operational criterion for identifying multifractal criticality beyond conventional approaches based solely on a single space.
An important open direction is to examine interacting many-body systems, where quasiperiodicity and interactions can produce many-body localized and critical phases. It will be particularly interesting to determine whether a similar scaling correspondence emerges between real-space and Fock-space or momentum-space structures of many-body eigenstates, and whether the Liu–Xia criterion can be extended to provide a precise characterization of many-body criticality.
@@ 1. Real-space Lyapunov exponent $\gamma$
For a given energy $E$, the real-space Lyapunov exponent $\gamma(E)$ characterizes the asymptotic exponential growth of the wave-function amplitude along the lattice. It is computed using the standard transfer-matrix method. Starting from an initial vector $\mathbf{v}_{0}$ (e.g., $(1,0)^{\mathrm{T}}$), one iterates the site-dependent transfer matrix $\mathbf{T}_{n}(E)$,
@@
After $N$ steps, the total transfer matrix is $\boldsymbol{\Pi}_{N}(E)=\mathbf{T}_{N-1}\cdots\mathbf{T}_{0}$, and the Lyapunov exponent is defined as
$\gamma(E)=\lim_{N\to\infty}\frac{1}{N}\ln\left\|\boldsymbol{\Pi}_{N}(E)\right\|.$ (10)
In numerical implementation, periodic renormalization is applied to prevent overflow, and $\gamma(E)$ is obtained from the averaged logarithmic growth rate for large $N$ (typically $N\sim 10^{5}$).
@@ 2. Momentum-space Lyapunov exponent $\gamma_{m}$
The momentum-space Lyapunov exponent $\gamma_{m}(E)$ is computed using an analogous transfer-matrix formulation in momentum space. After Fourier transforming the eigenvalue equation, the momentum-space amplitudes $\tilde{\psi}_{k}$ satisfy a linear recurrence relation. For quasiperiodic potentials with a single incommensurate frequency, this relation forms an effective one-dimensional chain,
@@
which can be rewritten as
$\begin{pmatrix}\tilde{\psi}_{k+Q}\\ \tilde{\psi}_{k}\end{pmatrix}=\mathbf{M}_{k}(E)\begin{pmatrix}\tilde{\psi}_{k}\\ \tilde{\psi}_{k-Q}\end{pmatrix}$. (12)
Iterating the corresponding transfer matrices yields
@@
where $\boldsymbol{\Pi}^{(m)}_{K}(E)$ denotes the product of momentum-space transfer matrices.
In both cases, the Lyapunov exponent quantifies the exponential growth rate of the transfer-matrix norm, with the only difference being whether the recurrence is defined in position or momentum space.
Appendix B Contrast Between Basis-Dependent IPR and Basis-Invariant Lyapunov Exponents
It is instructive to contrast the distinct roles of the Lyapunov exponent and the inverse participation ratio (IPR) in characterizing localization. The Lyapunov exponent is defined by the asymptotic exponential growth or decay of solutions to a linear recurrence relation and thus represents an intrinsic dynamical quantity associated with the transfer-matrix structure. Consequently, it is independent of the choice of basis and depends only on the underlying recursive dynamics in a given representation (real or momentum space).
In contrast, the IPR depends explicitly on the expansion coefficients of the wave function in a chosen basis and quantifies the degree of localization within that representation. Because higher-order moments of the wave-function amplitudes are not invariant under unitary transformations, the IPR is inherently basis dependent and generally differs between position and momentum space.
This distinction is central to the present work. While the simultaneous vanishing of Lyapunov exponents in dual spaces provides a basis-invariant criterion for criticality, the emergence of comparable scaling of the IPR in dual spaces constitutes a nontrivial physical signature. It indicates that, despite its basis dependence, the IPR reflects the reduced asymmetry between position and momentum space at criticality.
@@ Appendix C Fourier transformation for the QNE model
We show that the real-space equation
@@
transforms into the momentum-space equation
@@
under a discrete Fourier transform.
1. Fourier transform
Multiplying Eq. (14) by $\cos(2\pi\alpha n)$ gives
@@ 1. Fourier transform
2. Momentum-space equation
Using standard trigonometric identities, the Fourier transform yields
@@ 2. Momentum-space equation
Substituting into Eq. (16) gives
@@
which reduces to Eq. (15).